OpenStudy (anonymous):

61. Solve the equation 2a = 3a + 4c for a A. a = -4c B. a = 3 + 4c C. a = 4c – 5

OpenStudy (jagatuba):

Start by subtracting 3a from both sides.

OpenStudy (anonymous):

how?

OpenStudy (jagatuba):

2a=3a+4c -3a -3a -a = 4c Does that make sense?

OpenStudy (anonymous):

sort of

OpenStudy (jagatuba):

2a-3a=-a and (3a-3a)+4c = (0)+4c or just 4c

OpenStudy (jagatuba):

so now you have -a=4c so you want to multiply each side by -1 ro change the sign of a.

OpenStudy (anonymous):

-4c

OpenStudy (anonymous):

thanks for the help :D

OpenStudy (jagatuba):

correct

OpenStudy (jagatuba):

your welcome